# Robust Adaptive Discrete-Time Control Barrier Certificate

### v1

@@ 3.2 Robust adaptive control barrier certificate
The robust adapitve CBF (14) is built using the true parametric error $e_{\theta,t}$ instead of the set-induced error bound as in [25]. In fact, unlike what has been reported in [25], the invariance property of the safe sets defined by the maximized error bound-based robust adaptive CBFs can only be derived in discrete time, but not in continuous time. More details are given in Appendix B.
3.3 Practical implementation
In general, it is not possible to obtain an analytical expression of the safe policy $\pi_{\mathrm{ras},t}(\cdot;p)$. However, in practice, it suffices to only numerically compute a single safe input $u_{\mathrm{ras},t}(p)=\pi_{\mathrm{ras},t}(x_{t};p)$ by solving (19) for a specific $x_{t}$ at time step $t$. Efficient computation of $u_{\mathrm{ras},t}(p)$ requires fast characterization of $\mathcal{U}_{\mathrm{ras}}(x_{t};\hat{\theta}_{t},p)$ by establishing the inequality (16), which depends on $x_{t}$, $\hat{\theta}_{t}$, and $\varepsilon_{\theta,t}(p)$. However, computing $\varepsilon_{\theta,t}(2)$ requires solving a non-convex optimization problem. An alternative is using $\varepsilon_{\theta,t}(1)$, since then the optimization problem in (11b) can be reformulated as a linear program and solved efficiently [40]. However, since $\varepsilon_{\theta,t}(2)\leq\varepsilon_{\theta,t}(1)$, it follows that $\mathcal{U}_{\mathrm{ras}}(x;\hat{\theta}_{t},1)\subseteq\mathcal{U}_{\mathrm{ras}}(x;\hat{\theta}_{t},2)$ for all $x\in\mathcal{X}$, meaning the CBC (16) with $\varepsilon_{\theta,t}(1)$ is more conservative.
Remark 3.5.
@@ 3.3 Practical implementation
Employing $\varepsilon_{\theta,t}(1)$ in (16) results in a more conservative safe controller. Thus, $u_{\mathrm{ras},t}(1)$ is likely to be more suboptimal than $u_{\mathrm{ras},t}(2)$, especially in the context of optimal control when the policy $\pi_{\mathrm{nom},t}$ is (implicitly) given by methods using discrete-time control Lyapunov functions (CLFs) [6] or MPC [20], [21], [30].
In general, the optimization problem of the safety filter (19) is non-convex and cannot be solved efficiently. To address this issue, it is beneficial to construct a concave barrier candidate $B(\cdot)$ such that $\mathcal{U}_{\mathrm{ras}}(x;\hat{\theta}_{t},p)$ is convex, and thus computing $u_{\mathrm{ras},t}(p)$ requires solving a convex optimization problem, for which many efficient algorithms with convergence and optimality guarantees exist [7]. Note that concavity of $B(\cdot)$ is frequently assumed for DT-CBFs in the literature [2], [14], [47].
4 Online identification & Causal Synthesis
@@
In this section, following the adaptive control approach for linear systems [26], we instantiate an estimator $\mathtt{EST}$ using the least-squares-based recursive update with set-membership projection [19]. In addition, we present a causal integration of the safety filter and the estimator to guarantee that computing $u_{\mathrm{ras},t}(p)$ does not rely on any information from time step $t+1$ onward.
@@
For $t\geq 1$, both $\hat{\theta}_{t}$ and $\Theta_{t}$ are determined by the information flow $\mathcal{F}_{t-1}:=(x_{0},u_{0},x_{1},u_{1},\dots,x_{t-1})$, and a causal synthesis means computing $u_{t}$ based on $\mathcal{F}_{t}$. The initial estimate satisfies $\hat{\theta}_{0}\in\Theta_{0}$ with $\Theta_{0}=\Theta$. Given $x_{t}$, the residual is defined by $r_{t}:=x_{t}-f_{\mathrm{d}}(x_{t-1})-g(x_{t-1})u_{t-1}$, and it holds that $r_{t}+[\bm{\phi}(x_{t-1})]^{\top}\theta^{\ast}\in\mathcal{W}$, which means $\theta^{\ast}\in\Delta_{\theta,t}$ with the non-falsified set $\Delta_{\theta,t}$ given by
@@
$\Delta_{\theta,t}=\left\{\theta\in\mathbb{R}^{p}\mid H_{w}[\bm{\phi}(x_{t-1})]^{\top}\theta\leq\mathbf{1}-H_{w}r_{t}\right\}.\vskip-5.69046pt$ (21)
The set estimate is updated as
$\Theta_{t+1}=\Theta_{t}\cap\Delta_{\theta,t}.\vskip-5.69046pt$ (22)
For the point estimate, a prior is first computed based on the RLS estimation as
$\hat{\theta}_{t+1}^{[\mathrm{p}]}=\hat{\theta}_{t}-\mu_{t}\bm{\phi}(x_{t-1})\left(x_{t}-f(x_{t-1},u_{t-1};\hat{\theta}_{t})\right),\vskip-5.69046pt$ (23)
where $\mu_{t}\in(0,[\|\bm{\phi}(x_{t-1})\|^{2}]^{-1})$ is the learning rate. The final updated estimate $\hat{\theta}_{t+1}$ is computed by set-membership projection as
@@
and $\delta_{\theta,t}$ in (13) can be obtained accordingly. It is clear that $\delta_{\theta,t}$ is determined by $\mathcal{F}_{t}$, and therefore $u_{t}$, as computed as in (19), also depends on $\mathcal{F}_{t}$. In short, causality is ensured by synchronizing the estimates $\hat{\theta}_{t}$ and $\Theta_{t}$ with the one-step-delayed information flow $\mathcal{F}_{t-1}$.
@@
$x_{0}$, $\hat{\theta}_{0}$, $\Theta$, $\overline{w}$, $L_{B}$, $\Gamma$, $\{\pi_{\mathrm{nom},t}\}^{T-1}_{t=0}$
@@
  Determine $\mathcal{U}_{\mathrm{ras}}(x;\hat{\theta}_{t},p)$ $\triangleright$ (16)
@@
Algorithm 1 outlines the core steps of the proposed robust adaptive safe control with online estimation. The required inputs are the initial state $x_{0}$, initial estimates $\hat{\theta}_{0}$ and $\Theta$, constant parameters $\overline{w}$, $L_{B}$, and $\Gamma$, along with nominal policies $\{\pi_{\mathrm{nom},t}\}^{T-1}_{t=0}$. At time step $t$, $\hat{\theta}_{t}$ and $\delta_{\theta,t}$ are first computed to define $\mathcal{U}_{\mathrm{ras}}(x;\hat{\theta}_{t},p)$, and $\hat{\theta}_{t+1}$ and $\Theta_{t+1}$ will be used at $t+1$. Given $\mathcal{U}_{\mathrm{ras}}(x;\hat{\theta}_{t},p)$, a safe $u_{t}$ is then computed using the safety filter, and the system state evolves to $x_{t+1}$. It is also worth noting that, under Algorithm 1, we have $\Theta_{1}=\Theta_{0}$ and $\hat{\theta}_{1}=\hat{\theta}_{0}$, meaning the non-trivial update starts from $t=1$.
@@
The safe control scheme in Algorithm 1 also applies when the $1$-norm error bound $\varepsilon_{\theta,t}(1)$ is used. Moreover, given a model-based baseline policy $\pi_{\mathrm{nom}}$, the nominal policies $\{\pi_{\mathrm{nom},t}\}^{T-1}_{t=0}$ can also be generated by combining $\mathtt{EST}$, following the certainty-equivalence principle in adaptive control with online estimation [22], [26].
@@ 5 Inherent robustness
Inspired by [46], we discuss the inherent robustness property of the CBC in (16). Specifically, when $x_{t_{0}}$ is only safe but not robustly safe (i.e., $x_{t_{0}}\in\mathcal{S}\setminus\mathcal{S}_{\mathrm{ra},t_{0}}$) or even unsafe (i.e., $x_{t_{0}}\notin\mathcal{S}$) for some $t_{0}\in\mathbb{N}$ due to unexpected perturbations [46] or poor parameter estimates, inherent robustness ensures that the safe inputs derived from the safety filter (19) based on the CBC (16) are able to steer the state towards the sequence of robust adaptive safe sets $\{\mathcal{S}_{\mathrm{ra},t}\}^{\infty}_{t=t_{0}}$ asymptotically, i.e., $x_{t}\in\mathcal{S}_{\mathrm{ra},t}$ when $t\to\infty$. This asymptotic convergence w.r.t. a sequence of sets is formally given as follows:
@@
Assume that $\{\mathcal{X}_{t}\}^{\infty}_{t=0}$ with $\mathcal{X}_{t}\subseteq\overline{\mathcal{X}}$ is sequentially positively invariant (cf. Definition 3.1) for a general time-varying system $x_{t+1}=h_{t}(x_{t})+w_{t}$ for all $w_{t}\in\mathcal{W}$, where $h_{t}:\mathcal{X}\to\mathbb{R}^{n}$ is continuous for all $t\in\overline{\mathbb{N}}$. Then $x_{t}$ converges to $\{\mathcal{X}_{t}\}^{\infty}_{t=0}$ asymptotically on $\overline{\mathcal{X}}$ if $\lim_{t\to\infty}\mathrm{dist}_{\mathcal{X}_{t}}(x_{t})=0$ for all $x_{0}\in\overline{\mathcal{X}}$.
Following (15), we define the corresponding time-varying energy functions on $\mathcal{X}$ as follows:
@@
Let Assumption 1 hold. If $B:\mathcal{X}\to\mathbb{R}$ is a robust adaptive DT-CBF for (6) and $\mathtt{EST}$ for $T=\infty$ (cf. Definition 3.3), then $x_{t}$ converges to $\{\mathcal{S}_{\mathrm{ra},t}\}^{\infty}_{t=0}$ asymptotically on $\mathcal{X}$ (cf. Definition 5.1) for the system (6) with $u_{t}\in\mathcal{U}_{\mathrm{ras}}(x_{t};\hat{\theta}_{t},p)$.
@@ 6 Application Example
In this section, we consider the ACC problem, which is frequently used as a benchmark in safe control using CBFs [4], [24], [38], [45]. The continuous-time model of the ego vehicle is given by
@@
where $v$ and $d$ represent the velocity of the ego vehicle and the distance between the ego vehicle and the front vehicle, respectively. The parameter $M$ denotes the effective mass of the ego vehicle. The nominal velocity of the front vehicle, denoted by $v_{\mathrm{f}}$, is unknown but constrained within a known interval $v_{\text{lb}},v_{\text{ub}}]$. The control input $u$ represents the traction force applied by the ego vehicle, while the resistance force $F_{\mathrm{r}}(v)$ accounts for drag effects and is modeled as $F_{\mathrm{r}}(v)=F_{\text{roll}}+\mu_{\text{vis}}v+\mu_{\text{aero}}v^{2}$ for $v>0$, where $F_{\text{roll}}$ is the Coloumb friction force, $\mu_{\text{vis}}$ is the viscous force coefficient, and $\mu_{\text{aero}}$ is the aerodynamic drag coefficient [38], [45]. Both $F_{\text{roll}}$ and $\mu_{\text{vis}}$ are determined empirically offline [45], whereas $\mu_{\text{aero}}$ is identified online. The disturbance $w_{v}$ is included to capture velocity changes due to unmodeled resistance forces (e.g., wind effects), and $w_{d}$ characterizes stochastic variations of $v_{\mathrm{f}}$. These disturbances are bounded as $|w_{v}|\leq\overline{w}_{v}$ and $|w_{d}|\leq\overline{w}_{d}$, and the aerodynamic coefficient satisfies $\mu_{\text{aero}}\in[\mu^{-},\mu^{+}]$, where $\overline{w}_{v}$, $\overline{w}_{d}$, $\mu^{-}$, and $\mu^{+}$ are known constants. For simplicity and real-time computational efficiency [11], [37], forward Euler discretization is applied to obtain the discrete-time model as
@@
where $\Delta t=0.1[$\mathrm{s}$\$] is the sampling period. The discretized ACC model (27) is consistent with (2.2) with box-constrained $w_{t}=[w_{v,t}\Delta t,w_{d,t}\Delta t]^{\top}$ and $\theta=[\mu_{\text{aero}},v_{\mathrm{f}}]^{\top}$ such that (2) and (3) are satisfied. The safety constraint is $d\geq 1.8v$ [38], and the barrier function $B$ is chosen as B(v,d)=d1.8vaB(v,d)=d-1.8v-a, where $a\geq 0$ is user-specified safety margin. The nominal controller is an adaptive MPC controller equiped with $\mathtt{EST}$ using the certainty-equivalence principle [20], [22], [26], and the objective of the MPC is to track a reference velocity $v_{\text{ref}}=30[$\mathrm{m}\mathrm{/}\mathrm{s}$\$].
@@
Figure 1: $\mathtt{EST}$ estimation performance. The shaded area is the min-max envelope encompassing different disturbance realizations.
The simulations [^4] The full codebase and specific implementations are available at: https://github.com/lcrekko/ra_cbf_mpc. are conducted using Python 3.12. The performance of the online parameter estimation module $\mathtt{EST}$ is illustrated in Figure 1, which demonstrates that the true parameters are accurately recovered over time. The results related to safety are presented in Figure 2. As shown in the middle subplot, the proposed robust adaptive CBF-based safety filter ensures constraint satisfaction even in the presence of disturbances and model mismatch. This is achieved by actively modifying unsafe control inputs when safety is at risk, as evidenced in the bottom subplot of the same figure.
Based on Figure 2, the need for an explicit safety mechanism becomes apparent when comparing with the baseline adaptive MPC controller, which fails to enforce the safety constraint. Although the robust CBF controller also maintains safety, it reacts more slowly to critical situations compared to the robust adaptive CBF approach. As a result, from approximately $7$ seconds onward, stronger braking actions–i.e., more conservative control inputs–are required to uphold safety.
@@
Figure 2: Comparison of adaptive MPC with robust adaptive CBF-based safety filter (aMPC-raCBF), MPC with worst-case robust CBF-based safety filter (MPC-rCBF), and adaptive MPC (aMPC). The shaded area is the min-max envelope encompassing different disturbance realizations. Zoomed-in subplot is provided for each subfigure.
7 Conclusions

### v2

@@ 3.3 Practical implementation
Employing $\varepsilon_{\theta,t}(1)$ in (17) results in a more conservative safe controller. Thus, $u_{\mathrm{ras},t}(1)$ is likely to be more suboptimal than $u_{\mathrm{ras},t}(2)$, especially in the context of optimal control when the policy $\pi_{\mathrm{nom},t}$ is (implicitly) given by methods using discrete-time control Lyapunov functions (CLFs) [5], reinforcement learning [9], or approximate MPC [13], [20]. An alternative approach is to use the $\infty$-norm uniformly when quantifying all the error bounds (i.e., (5), (9b), (12), and (16)) and the Lipschitz condition in Assumption 1.
4 Online estimation & Causal Synthesis
@@
In this section, following the adaptive control approach for linear systems [26], we instantiate an estimator $\mathtt{EST}$ using the clipped recursive-least-square (RLS) update with set-membership projection. In addition, we present a causal integration of the safety filter and the estimator to guarantee that computing $u_{\mathrm{ras},t}(p)$ does not rely on any information from time step $t+1$ onward.
@@
For $t\geq 1$, both $\hat{\theta}_{t}$ and $\Theta_{t}$ are determined by the information flow $\mathcal{F}_{t-1}:=(x_{0},u_{0},x_{1},u_{1},\dots,x_{t-1})$, and a causal synthesis means computing $u_{t}$ based on $\mathcal{F}_{t}$. The initial estimate satisfies $\hat{\theta}_{0}\in\Theta_{0}$ with $\Theta_{0}=\Theta$. Given $x_{t}$, the residual is defined by $r_{t}:=x_{t}-f_{\mathrm{d}}(x_{t-1})-g(x_{t-1})u_{t-1}$, and it holds that $r_{t}+[\bm{\phi}(x_{t-1})]^{\top}\theta^{\ast}\in\mathcal{W}$, i.e., $H_{w}(r_{t}+[\bm{\phi}(x_{t-1})]^{\top}\theta^{\ast})\leq\mathbf{1}_{n_{w}}$ by (2). Equivalently, $\theta^{\ast}\in\Delta_{\theta,t}$ with the non-falsified polytopic set $\Delta_{\theta,t}$ given by
@@
$\Delta_{\theta,t}=\left\{\theta\in\mathbb{R}^{q}\mid H_{w}[\bm{\phi}(x_{t-1})]^{\top}\theta\leq\mathbf{1}_{n_{w}}\hskip-2.27626pt-\hskip-2.27626ptH_{w}r_{t}\right\}.\vskip-5.69046pt$ (21)
Given the current set estimate $\Theta_{t}=\{\theta\in\mathbb{R}^{q}\mid H_{\theta,t}\theta\leq h_{\theta,t}\}$, the updated set estimate is computed as
$\Theta_{t+1}=\Theta_{t}\cap\Delta_{\theta,t}\vskip-5.69046pt$ (22)
which is a polytope described by an augmented set of hyperplanes [^5] Unbounded growth in the number of hyperplanes can be avoided by overapproximating the intersection with a hyperrectangle (or hyberbox) of fixed dimension [26].contributing from $\Theta_{t}$ and $\Delta_{\theta,t}$. For the point estimate, a prior is first computed based on the clipped RLS parameter update as
$\hat{\theta}_{t+1}^{[\mathrm{p}]}=\hat{\theta}_{t}-\mu_{t}\underbrace{\bm{\phi}(x_{t-1})\left(x_{t}-f(x_{t-1},u_{t-1};\hat{\theta}_{t})\right)}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}:=\psi_{\theta,t}}},\vskip-5.69046pt$ (23)
where $\mu_{t}=\min\left\{(\left\|\psi_{\theta,t}\right\|+\epsilon)^{-1},\|\bm{\phi}(x_{t-1})\|^{-2}\right\}$ is the clipped learning rate with $\epsilon>0$ being a small constant. Note that (23) degenerates to a gradient descent update when the learning rate is constant [26], and the proposed clipped RLS update follows gradient clipping [16]. The final updated estimate $\hat{\theta}_{t+1}$ is computed by set-membership projection as
@@
and $\delta_{\theta,t}$ in (11) can be obtained accordingly. Due to $\|\mu_{t}\psi_{\theta,t}\|\leq 1$ and the triangle inequality, it holds that $\|\delta_{\theta,t}\|\leq 1$. Since $\delta_{\theta,t}$ is determined by $\mathcal{F}_{t}$, and thus $u_{t}$, as computed as in (20), also depends on $\mathcal{F}_{t}$. In short, causality is ensured by synchronizing the estimates $\hat{\theta}_{t}$ and $\Theta_{t}$ with the one-step-delayed information flow $\mathcal{F}_{t-1}$.
Algorithm 1 outlines the core steps of the proposed safe control approach. The required inputs are the CBF $B$ and its associated $\alpha\in\mathcal{K}_{\mathrm{e},\infty}\cap\mathcal{C}_{\mathrm{SL}}$, initial variables $x_{0}$, $\hat{\theta}_{0}$ and $\Theta$, and constants $\overline{w}$, $L_{B,x}$, $L_{B,\theta}$, and $\Gamma$. At time step $t$, $\hat{\theta}_{t}$ and $\delta_{\theta,t}$ are first computed to define $\mathcal{U}_{\mathrm{ras},t}(x_{t};p)$, and $\hat{\theta}_{t+1}$ and $\Theta_{t+1}$ will be used at $t+1$. Given $\mathcal{U}_{\mathrm{ras},t}(x;p)$, a safe $u_{t}$ is then computed using the safety filter, and the system state evolves to $x_{t+1}$. It is also worth noting that $\Theta_{1}=\Theta_{0}$ and $\hat{\theta}_{1}=\hat{\theta}_{0}$, meaning the non-trivial update starts from $t=1$.
@@
$B$, $\alpha$, $x_{0}$, $\hat{\theta}_{0}$, $\Theta$, $\overline{w}$, $L_{B,x}$, $L_{B,\theta}$, $\Gamma$
@@
  Determine $\mathcal{U}_{\mathrm{ras},t}(x;p)$ $\triangleright$ (15)
@@
The nominal policies $\{\pi_{\mathrm{nom},t}\}^{T-1}_{t=0}$ can also be designed by combining the estimator $\mathtt{EST}$ if model-based approaches are adopted, following the certainty-equivalence principle [21], [22] in adaptive control with online estimation [23], [26].
Remark 4.2.
Convergence of the estimate $\hat{\theta}_{t}$ to the true parameter $\theta^{\ast}$ typically requires additional conditions such as persistent excitation [24], [29], which can be achieved through active exploration (e.g., injecting noise into the input [9]). In contrast, the proposed safe control approach relies only on the uncertainty set $\Theta_{t}$ that contains the true parameter $\theta^{\ast}$, such that safety is ensured at all times by robustification against $\Theta_{t}$. Although convergence allows conservatism to be reduced maximally as $E_{\theta,t}(x)$ in (17) approaches zero, the safety guarantees developed in this work do not rely on achieving convergence. The explicit design of exploration strategies to ensure convergence is therefore beyond the scope of the current paper.
@@ 5 Inherent robustness
Inspired by [43], we discuss the inherent robustness property of the input condition in (17). Specifically, when $x_{t_{0}}$ is only safe but not robustly safe (i.e., $x_{t_{0}}\in\mathcal{S}\setminus\mathcal{S}_{\mathrm{ra},t_{0}}$) or even unsafe (i.e., $x_{t_{0}}\notin\mathcal{S}$) for some $t_{0}\in\mathbb{N}$ due to unexpected perturbations [43] or poor parameter estimates, inherent robustness ensures that the safe inputs derived from the safety filter (20) based on the condition (17) are able to steer the state towards the sequence of robust adaptive safe sets $\{\mathcal{S}_{\mathrm{ra},t}\}^{\infty}_{t=t_{0}}$ asymptotically, i.e., $x_{t}\in\mathcal{S}_{\mathrm{ra},t}$ when $t\to\infty$. This asymptotic convergence w.r.t. a sequence of sets is formally defined as follows:
@@
Assume that $\{\mathcal{X}_{t}\}^{\infty}_{t=0}$ with $\mathcal{X}_{t}\subseteq\overline{\mathcal{X}}$ is sequentially positively invariant (cf. Definition 3.1) for a general time-varying system $x_{t+1}=h_{t}(x_{t})+w_{t}$ for all $w_{t}\in\mathcal{W}$, where $h_{t}:\mathcal{X}\to\mathbb{R}^{n}$ is continuous for all $t\in\overline{\mathbb{N}}$. Then $x_{t}$ converges to $\{\mathcal{X}_{t}\}^{\infty}_{t=0}$ asymptotically on $\overline{\mathcal{X}}$ is defined as $\lim_{t\to\infty}\mathrm{dist}_{\mathcal{X}_{t}}(x_{t})=0$ for all $x_{0}\in\overline{\mathcal{X}}$.
Following (14), we define the corresponding time-varying continuous energy functions on $\mathcal{X}$ as follows:
@@
Let Assumption 1 hold. If $B:\mathcal{X}\times\Theta\to\mathbb{R}$ is a robust adaptive DT-CBF for (6) and $\mathtt{EST}$ (cf. Definition 3.2), then $x_{t}$ converges to $\{\mathcal{S}_{\mathrm{ra},t}\}^{\infty}_{t=0}$ asymptotically on $\mathcal{X}$ (cf. Definition 5.1) for the system (6) with $u_{t}\in\mathcal{U}_{\mathrm{ras},t}(x_{t};p)$.
@@ 6 Application Example
Figure 2: Diagram of the proposed safe control approach for the adaptive cruise control problem using robust adaptive discrete-time control barrier functions.
In this section, we apply the proposed DT-CBF-based safe control to an adaptive cruise control (ACC) problem, a standard benchmark in safe control [3], [39], [42]. The ego vehicle dynamics are
@@
where $v$, $M$, and $d$ denote velocity, vehicle mass, and the distance to the front vehicle, respectively. The nominal velocity of the front vehicle $v_{\mathrm{f}}\in[v_{\text{lb}},v_{\text{ub}}]$ is unknown. The input $u\in[-u_{\max},u_{\max}]$ represents the traction force applied by the ego vehicle. The resistance force $F_{\mathrm{r}}(v)$ accounts for drag effects and is modeled as $F_{\mathrm{r}}(v)=F_{\text{roll}}+\mu_{\text{vis}}v+\mu_{\text{aero}}v^{2}$ for $v>0$, where $F_{\text{roll}}$ is the Coloumb friction force, $\mu_{\text{vis}}$ is the viscous force coefficient, and $\mu_{\text{aero}}$ is the aerodynamic drag coefficient [39], [42]. Both $F_{\text{roll}}$ and $\mu_{\text{vis}}$ are determined empirically offline [42], whereas $\mu_{\text{aero}}\in[\mu^{-},\mu^{+}]$ is identified online. The disturbance $w_{v}\in[-\overline{w}_{v},\overline{w}_{v}]$ is included to capture wind effects, and $w_{d}\in[-\overline{w}_{d},\overline{w}_{d}]$ characterizes variations of $v_{\mathrm{f}}$. Table 1 summarizes the nominal parameters and the known bounds. For simplicity and computational efficiency [11], [38], Euler discretization with sampling period $\Delta t=0.1[$\mathrm{s}$\$] is applied to obtain
[…]
The model (27) is of the general form in (2.2) with box-constrained $w_{t}=[w_{v,t}\Delta t,w_{d,t}\Delta t]^{\top}$ and $\theta=[\mu_{\text{aero}},v_{\mathrm{f}}]^{\top}$ satisfying (2) and (3). Safety is enforced through the time-headway constraint $d\geq 1.8v$ [39], and the CBF [^6] For the considered adaptive cruise control example, we adopt the parameter-free CBF as in [39].is chosen as B(v,d)=d1.8v0.5B(v,d)=d-1.8v-0.5 with $\alpha(x)=(1-10^{-4})x$. It can be numerically verified that the CBC in (15) is satisfied for $p=2$ and $\Gamma=10^{5}\mathbf{I}_{2}$. The nominal controller is an unconstrained myopic (i.e., prediction horizon $N=1$) adaptive MPC (aMPC) controller [^7] Our contribution is not centered on MPC; it is used only as a nominal controller for numerical illustration. tracking $v_{\text{ref}}=30[\text{m/s}]$, equiped with the estimator that computes $\hat{\mu}_{\text{aero}}$ and $\hat{v}_{\text{f}}$ as in Section 4. The nominal input from aMPC is filtered by the safety filter in (20) to obtain the safe input. Fig. 2 presents a diagram that highlights the connections between the theoretical components and their implementation.
@@
Figure 3: $\mathtt{EST}$ performance. The shaded area is the min-max envelope encompassing different disturbance realizations.
@@
Figure 4: Comparison of adaptive MPC with robust adaptive CBF-based safety filter (aMPC-raCBF), MPC with worst-case robust CBF-based safety filter (MPC-rCBF), adaptive MPC (aMPC), and robust adaptive MPC (raMPC) [18]. The shaded area is the min-max envelope encompassing different disturbance realizations.
Table 1: Parameters and bounds used in the numerical simulation.
Nominal parameters Bounds
Parameter Value Units Parameter Value Units
$M$ $1650$ $\mathrm{kg}$ $v_{\text{lb}},v_{\text{ub}}] [$2020, $32$] $\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$
$F_{\text{roll}}$ $125$ $\mathrm{kg}\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-2}$ $\mu^{-},\mu^{+}] [$$0.1$, $0.6$] $\mathrm{kg}\text{\,}{\mathrm{m}}^{-1}$
$\mu_{\text{vis}}$ $1.2$ $\mathrm{kg}\text{\,}{\mathrm{s}}^{-1}$ $\overline{w}_{v}$ 22 $\mathrm{m}\text{\,}{\mathrm{s}}^{-2}$
$\mu^{\ast}_{\text{aero}}$ $0.55$ $\mathrm{kg}\text{\,}{\mathrm{m}}^{-1}$ $\overline{w}_{d}$ 55 $\mathrm{m}\text{\,}{\mathrm{s}}^{-2}$
$v^{\ast}_{\text{f}}$ $22$ $\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$ $u_{\max}$ 10 00010\,000 $\mathrm{kg}\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-2}$
The simulations [^8] The full codebase and specific implementations are available at: https://github.com/lcrekko/ra_cbf_mpc. are conducted using Python 3.12. The performance of the online estimator $\mathtt{EST}$ is illustrated in Fig. 3, which demonstrates that the true parameters are accurately recovered over time. The results related to safety are presented in Fig. 4. As shown in the middle subplot, the proposed robust adaptive CBF-based safety filter ensures constraint satisfaction even in the presence of disturbances and model mismatch. This is achieved by actively modifying unsafe control inputs when safety is at risk, as evidenced in the bottom subplot of the same figure. Based on Fig. 4, the need for an explicit safety mechanism becomes apparent when comparing with the baseline unconstrained aMPC controller, which fails to enforce the safety constraint. Although method using robust CBF and robust adaptive MPC with constraint tightening [18] also maintains safety, they react more slowly to critical situations compared to the safe control approach using robust adaptive CBFs. As a result, from approximately $7$ seconds onward, stronger braking actions–i.e., more conservative control inputs–are required to uphold safety. Although robust adaptive MPC [17], [18], [27] provides an alternative approach to safe control under uncertainty, methods based on CBFs are modular by design and can be combined with nominal controllers that do not inherently guarantee safety, e.g., reinforcement learning [8] and approximate MPC [13], [20].
7 Conclusions & Future work